OpenStudy (michellemn):

Match each shape with its surface area. Use three point one four for pi and round to the nearest whole number where necessary. Match Term 1. square pyramid: base length 5 meters, slant height 10 meters 2. cone: radius 6 meters, slant height 8 meters 3. square pyramid: base length 6 meters, slant height eleven meters 4. cone: radius 6 meters, slant height 5 meters Definition A) one hundred twenty-five square meters B) two hundred sixty-four square meters C) one hundred sixty-eight square meters D) two hundred seven square meters

OpenStudy (michellemn):

Match the term with the definition. Can someone help please?!

OpenStudy (anonymous):

i will help i think

OpenStudy (anonymous):

FOR CONE : Lateral Surface area = π×r×l ; l = slant height FOR Right Square Pyramid it is : =a2+2aa24+h−−−−−−√ a= base edge h - height (NOT SLANT HEIGHT) If you are given with slant height in case of Right Square Pyramid then : SA = a2+2×a×l

OpenStudy (michellemn):

@iGreen

OpenStudy (igreen):

Formulas: Cone: \(SA = \pi r s + \pi r^2\) Square Pyramid: \(SA = \pi r^2 + \pi r^2 s\)

OpenStudy (michellemn):

Thanks that's all I needed!

OpenStudy (igreen):

Oh, really? No problem.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @michellemn

OpenStudy (anonymous):

oh ok so what i gave u did not work right?

OpenStudy (michellemn):

no its just that i dont understand the �

OpenStudy (michellemn):

Thats basically all i see in your post

OpenStudy (anonymous):

is not there anymore go up and see

OpenStudy (michellemn):

um its still there

OpenStudy (michellemn):

ok its gone now

OpenStudy (michellemn):

thanks both of you!!

OpenStudy (igreen):

@michellemn You have to click on a different post and click back to this one. Or you can reload the page..

OpenStudy (anonymous):

see